OpenStudy (idealist10):

Find a particular solution of y"+2y'+2y=e^(-x)(8cosx-6sinx).

OpenStudy (idealist10):

\[y _{p}=xe ^{-x}(Acosx+Bsinx)\]

OpenStudy (idealist10):

Is my initial guess right above?

OpenStudy (irishboy123):

you've taken the complementary solution and just tines'd it by x. doesn't work that way. you could, having worked the complementary solution, then go through the Variation of Params route.

OpenStudy (irishboy123):

to be sure, there are other ways to go about it eg you can re-write the RHS as \(10 e^{-x} (\frac{4}{5} \cos x - \frac{3}{5} \sin x ) = 10 e^{-x} \mathcal{Re } \{ e^{i(x + \alpha)}\}\) and work it like that, ie solve \(z_p = x_p + i y_p\), but looks like an awful mess or you can guess the solution, aka Method of Undetermined Coefficients, but I have no idea OTTOMF what a good guess would be Var of Params should just work, described here: http://tutorial.math.lamar.edu/Classes/DE/VariationofParameters.aspx pretty sure someone else will think of other ways too :-)

OpenStudy (thomas5267):

\(e^{-x}(A\cos(x)-B\sin(x))\) looks like a good guess to me.

hartnn (hartnn):

\(\color{blue}{\text{Originally Posted by}}\) @Idealist10 \[y _{p}=xe ^{-x}(Acosx+Bsinx)\] \(\color{blue}{\text{End of Quote}}\) i'd say the same thing as them, \[y _{p}=e ^{-x}(Acosx+Bsinx)\] is better

OpenStudy (idealist10):

\[y _{p}=e ^{-x}(Acosx+Bsinx)\] doesn't work. I tried.

hartnn (hartnn):

then you might have tried \[y _{p}=xe ^{-x}(Acosx+Bsinx)\] too? that should work ! if not, then even idk :P

OpenStudy (thomas5267):

The guess \(y_p=xe^{-x}\left(A\cos(x)+B\sin(x)\right)\) works. Try again.